# Advances in the Shannon Capacity of Graphs

### v1

@@ Abstract
This paper studies several research directions concerning the Shannon capacity of graphs. Building on Schrijver’s recent framework, we establish sufficient conditions under which the Shannon capacity of a polynomial in graphs equals the corresponding polynomial of the individual capacities, thereby simplifying their evaluation. We derive exact values and new bounds for the Shannon capacity of two families of graphs: the $q$-Kneser graphs and the Tadpole graphs. Furthermore, we construct graphs whose Shannon capacity is never attained by the independence number of any finite power of these graphs, including a countably infinite family of connected graphs with this property. We further prove an inequality relating the Shannon capacities of the strong product of graphs and their disjoint union, leading to streamlined proofs of known bounds.
@@ 1 Introduction
The Shannon capacity of graphs, introduced by Claude Shannon in [1], equals the maximum transmission rate at which a receiver can accurately recover information without error, with the communication channel represented as a graph. In this graph, the vertices represent the input symbols and any two vertices are adjacent if and only if the corresponding input symbols can be confused by the channel with some positive probability. This concept establishes a significant link among zero-error problems in information theory and graph theory, prominently featured in multiple surveys [2, 3, 4, 5, 6].
@@
While investigating the properties of the Shannon capacity, it is worth keeping in mind the more general concept of the asymptotic spectrum of graphs, introduced by Zuiddam [21, 22, 20, 19], which delineates a space of graph parameters that remain invariant under graph isomorphisms. This space is characterized by the following unique properties: additivity under disjoint union of graphs, multiplicativity under strong product of graphs, normalization for a simple graph with a single vertex, and monotonicity under graph complement homomorphisms. Building upon Strassen’s theory of the asymptotic spectra [23], a novel dual characterization of the Shannon capacity of a graph is derived in [21], expressed as the minimum over the elements of its asymptotic spectrum. By confirming that various graph invariants, including the Lovász $\vartheta$-function [11] and the fractional Haemers bound [13], are elements of the asymptotic spectrum of a graph (spectral points), it can be deduced that these elements indeed serve as upper bounds on the Shannon capacity of a graph. For further exploration, the comprehensive paper by Wigderson and Zuiddam [22] provides a survey on Strassen’s theory of the asymptotic spectra and its application areas, including the Shannon capacity of graphs.
Several properties of the Shannon capacity have been explored, and many questions are left open. One of the most basic open questions is the determination of the Shannon capacity of cycle graphs of odd length larger than 5.
@@
  • Section 5 determines sufficient conditions for the unattainability of the Shannon capacity by the independence number of any finite strong power of a graph. It first presents in an alternative streamlined way an approach by Guo and Watanabe [10]. It then introduces two other original approaches. One of the novelties in this section is the construction of an infinite family of connected graphs whose capacity is unattainable by any finite strong power of these graphs.
  • Section 6 derives the Shannon capacity of the q-Kneser graphs, using a generalized result of Erdos-Ko-Rado [25], and a result on the spectrum of the $q$-Kneser graphs [26]. This broadens the class of graphs for which the Shannon capacity is explicitly known.
  • Section 7 introduces a new inequality relating the Shannon capacity of the strong power of graphs to that of their disjoint union, and itidentifies several conditions under which equality holds. As an illustration of the applicability of this inequality, this section also presents an alternative shorter proof of a lower bound for the disjoint union of a graph with its complement, originally obtained by N. Alon [7], along with analogous bounds on the Lovász $\vartheta$-function of graphs.
@@ 2.1.1 Terminology
  • A walk in a graph $\mathsf{G}$ is a sequence of vertices in $\mathsf{G}$, where every two consecutive vertices in the sequence are adjacent in $\mathsf{G}$.
  • A trail in a graph is a walk with no repeated edges.
  • A path in a graph is a walk with no repeated vertices; consequently, a path has no repeated edges, so every path is a trail but a trail is not necessarily a path.
  • A cycle $\operatorname{\mathsf{C}}$ in a graph $\mathsf{G}$ is obtained by adding an edge to a path $\operatorname{\mathsf{P}}$ such that it gives a closed walk.
@@
  • An $r$-partite graph is a graph whose vertex set is a disjoint union of $r$ subsets such that no two vertices in the same subset are adjacent. If $r=2$, then $\mathsf{G}$ is a bipartite graph.
@@
  • A path graph on $n$ vertices is denoted by $\operatorname{\mathsf{P}}_{n}$, and its size is equal to n1n-1.
  • A cycle graph on $n$ vertices is called an $n$-cycle, and it is denoted by $\operatorname{\mathsf{C}}_{n}$ with an integer $n\geq 3$. The order and size of $\operatorname{\mathsf{C}}_{n}$ are equal to $n$, and $\operatorname{\mathsf{C}}_{n}$ is a bipartite graph if and only if $n\geq 4$ is even.
@@
Throughout this paper, the graphs under consideration are finite, simple, and undirected. The standard notation $n]\triangleq\{1,\ldots,n\}$, for every $n\in\mathbb{N}$, is also used.
@@ 2.1.2 Graph operations
Define the $k$-fold strong power of $\mathsf{G}$ as
$\mathsf{G}^{k}\triangleq\underbrace{\mathsf{G}\boxtimes\ldots\boxtimes\mathsf{G}.}_{k-1\text{ strong products}}$ (1)
Throughout this article, we will use another graph operation.
@@
Let $\mathsf{G}$ and $\mathsf{H}$ be two simple graphs. The disjoint union $\mathsf{G}+\mathsf{H}$ is a graph whose vertices set is $\operatorname{\mathsf{V}}(\mathsf{G})\cup\operatorname{\mathsf{V}}(\mathsf{H})$, and the edges set is $\operatorname{\mathsf{E}}(\mathsf{G})\cup\operatorname{\mathsf{E}}(\mathsf{H})$.
Let $m\in\mathbb{N}$, throughout this paper we use the following notation:
$m\mathsf{G}\triangleq\underbrace{\mathsf{G}+\ldots+\mathsf{G}.}_{m-1\text{ disjoint unions}}$ (2)
@@ 2.1.4 Fractional invariants of graphs
To properly define the four basic fractional invariants of a graph, we need to define the following four sets of functions.
@@
Let $\mathsf{G}$ be a simple graph. Define the four following sets of functions:
  • Define $\mathcal{F}_{I}(\mathsf{G})$ as the set of non-negative functions, $f\colon\operatorname{\mathsf{V}}(\mathsf{G})\to\mathbb{R}$, such that for every $\mathcal{I}\in\operatorname{\mathcal{I}}(\mathsf{G})$,
@@
  • Define $\mathcal{F}_{C}(\mathsf{G})$ as the set of non-negative functions, $f\colon\operatorname{\mathsf{V}}(\mathsf{G})\to\mathbb{R}$, such that for every $\mathcal{C}\in\operatorname{\mathcal{C}}(\mathsf{G})$,
@@
  • Define $\mathcal{G}_{I}(\mathsf{G})$ as the set of non-negative functions, $g\colon\operatorname{\mathcal{I}}(\mathsf{G})\to\mathbb{R}$, such that for every $v\in\operatorname{\mathsf{V}}(\mathsf{G})$,
    $\sum_{\mathcal{I}\in\operatorname{\mathcal{I}}(\mathsf{G}):v\in\mathcal{I}}g(\mathcal{I})\geq 1.$ (11)
  • Define $\mathcal{G}_{C}(\mathsf{G})$ as the set of non-negative functions, $g\colon\operatorname{\mathcal{C}}(\mathsf{G})\to\mathbb{R}$, such that for every $v\in\operatorname{\mathsf{V}}(\mathsf{G})$,
    $\sum_{\mathcal{C}\in\operatorname{\mathcal{C}}(\mathsf{G}):v\in\mathcal{C}}g(\mathcal{C})\geq 1.$ (12)
Four basic fractional invariants are next defined by linear programming.
@@
Using strong duality of linear programming we get the following important theorem.
@@
Let $\mathsf{G}$ be a simple graph. Then
@@
So, from now on we will use mostly the fractional independence number and the fractional chromatic number. And, when we calculate them we will use both of their equivalent linear programming representations.
Another useful property, that is obvious from Definition 2.13, is provided next.
@@
For a simple graph $\mathsf{G}$, the following holds:
@@
Finally, some properties of the fractional independence number are presented.
@@ 2.1.6 Some structured families of graphs
A corollary by Hales from 1973 regarding the connection between graph universality and the fractional independence number is presented next.
@@
A graph $\mathsf{G}$ is strongly regular with parameters $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ if it suffices:
@@
  • Two vertices $a$ and $b$ are adjacent if and only if $a-b\in(\mathbb{F}_{q}^{\times})^{2}$.
@@
[29] Let $\mathsf{G}=P(q)$ be a Paley graph with $q\equiv 1\text{ mod 4}$ a prime power. Then
@@ 2.2 The Shannon capacity of graphs
The concept of the Shannon capacity of a graph $\mathsf{G}$ was introduced by Claude E. Shannon in [1] to consider the largest information rate that can be achieved with zero-error communication. A discrete channel consists of
  • A finite input set $\mathcal{X}$.
  • A (possibly infinite) output set $\mathcal{Y}$.
  • A non-empty fan-out set $\mathcal{S}_{x}\subseteq\mathcal{Y}$ for every $x\in\mathcal{X}$.
In each channel use, a sender transmits an input $x\in\mathcal{X}$ and a receiver receives an arbitrary output in $\mathcal{S}_{x}$. It is possible to represent a DMC (discrete memoryless channel) by a confusion graph $\mathsf{G}$, which will be defined as follows
  • $\operatorname{\mathsf{V}}(\mathsf{G})=\mathcal{X}$ represents the symbols of the input alphabet to that channel.
  • $\operatorname{\mathsf{E}}(\mathsf{G})$ is the edge set of $\mathsf{G}$, where two distinct vertices in $\mathsf{G}$ are adjacent if the corresponding two input symbols from $\mathcal{X}$ are not distinguishable by the channel, i.e., they can produce an identical output symbol with some positive probability. Formally,
@@
Thus, the largest number of inputs a channel can communicate without error in a single use is the independence number $\alpha(\mathsf{G})$.
In the considered setting, the sender and the receiver agree in advance on an independent set $\mathcal{I}$ of a maximum size $\alpha(\mathsf{G})$, the sender transmits only inputs in $\mathcal{I}$, every received output is in the fan-out set of exactly one input in $\mathcal{I}$, and the receiver can correctly determine the transmitted input. Next, consider a transmission of $k$-length strings over a channel, where the channel is used $k\geq 1$ times, the sender transmits a sequence $x_{1}\ldots x_{k}$, and the receiver gets a sequence $y_{1}\ldots y_{k}$ of outputs, where $y_{i}\in\mathcal{S}_{x_{i}}$ for all $i\in[k]$. In this setup, $k$ uses of the channel are viewed as a single use of a super-channel:
  • Its input set is $\mathcal{X}^{k}$, and its output set is $\mathcal{Y}^{k}$.
  • The fan-out set of $(x_{1},\ldots,x_{k})\in\mathcal{X}^{k}$ is the cartesian product
    $\mathcal{S}_{x_{1}}\times\ldots\times\mathcal{S}_{x_{k}}.$ (25)
Note that two sequences $(x_{1},\ldots,x_{k}),(x_{1}^{\prime},\ldots,x_{k}^{\prime})\in\mathcal{X}^{k}$ are distinguishable by the channel if and only if there exists $i\in[k]$ such that
$S_{x_{i}}\cap S_{x_{i}^{\prime}}=\varnothing.$ (26)
Thus, it is possible to represent the larger channel by the strong powers of the confusion graph, and the $k$-th confusion graph is defined as thr $k$-fold strong power of $\mathsf{G}$:
Using the $k$-th confusion graph, the largest amount of information we can send through the channel with $k$-uses is the independence number of the $k$-th confusion graph, $\alpha(\mathsf{G}^{k})$. Thus, the maximum information rate per symbol that is achievable by using input strings of length $k$ is equal to
$\frac{1}{k}\log\alpha(\mathsf{G}^{k})=\log\sqrt[k]{\alpha(\mathsf{G}^{k})},\;\;\;k\in\mathbb{N}.$
And, by omitting the logarithm (as an increasing function) and supremizing over $k$, the Shannon capacity is defined as follows.
@@
$\Theta(\mathsf{G})\triangleq\sup_{k\in\mathbb{N}}\sqrt[k]{\alpha(\mathsf{G}^{k})}=\lim_{k\to\infty}\sqrt[k]{\alpha(\mathsf{G}^{k})}.$ (28)
Remark 2.1.
The righthand equality in (28) is derived from Theorem 2.1 and Fekete’s Lemma.
Now we provide several basic properties of the Shannon capacity which will be used throughout this paper.
@@
Let $\mathsf{G}$ be a simple graph and let $\ell\in\mathbb{N}$. Then
$\Theta(\mathsf{G}^{\ell})=\Theta(\mathsf{G})^{\ell}.$ (29)
@@
Let $\mathsf{G}$ be a simple graph and let $m\in\mathbb{N}$, then
$\Theta(m\mathsf{G})=m\Theta(\mathsf{G}).$ (30)
Theorem 2.10.
Let $\mathsf{G}_{1}$ and $\mathsf{G}_{2}$ be simple graphs. Then
@@
[1] Let $\mathsf{G}_{1}$ and $\mathsf{G}_{2}$ be simple graphs. Then
@@
For an elegant proof of Shannon’s theorem, see [24].
The following theorem is very important and central in this paper, thus, its proof is presented as well.
@@
[24] Let $\mathsf{G}$ and $\mathsf{H}$ be simple graphs. Then
@@
By Theorems 2.10 and 2.11, the claim is equivalent to
$\Theta(\mathsf{G}+\mathsf{H})>\Theta(\mathsf{G})+\Theta(\mathsf{H})\iff\Theta(\mathsf{G}\boxtimes\mathsf{H})>\Theta(\mathsf{G})\,\Theta(\mathsf{H}).$ (34)
We next prove both directions of the equivalence in (34).
  1. 1.
    Assume that $\Theta(\mathsf{G}\boxtimes\mathsf{H})>\Theta(\mathsf{G})\,\Theta(\mathsf{H})$. By Theorem 2.8,
    $\Theta(\mathsf{G}+\mathsf{H})^{2}$ $=\Theta((\mathsf{G}+\mathsf{H})^{2})$
    $=\Theta(\mathsf{G}^{2}+2\mathsf{G}\boxtimes\mathsf{H}+\mathsf{H}^{2}),$
    and by Theorems 2.10 and 2.11,
    $\Theta(\mathsf{G}^{2}+2\mathsf{G}\boxtimes\mathsf{H}+\mathsf{H}^{2})\geq\Theta(\mathsf{G})^{2}+2\Theta(\mathsf{G}\boxtimes\mathsf{H})+\Theta(\mathsf{H})^{2}.$
    Using the above assumption, we get
    $\Theta(\mathsf{G})^{2}+2\Theta(\mathsf{G}\boxtimes\mathsf{H})+\Theta(\mathsf{H})^{2}$ $>\Theta(\mathsf{G})^{2}+2\Theta(\mathsf{G})\,\Theta(\mathsf{H})+\Theta(\mathsf{H})^{2}$
    $=(\Theta(\mathsf{G})+\Theta(\mathsf{H}))^{2},$
    which gives
    $\Theta(\mathsf{G}+\mathsf{H})>\Theta(\mathsf{G})+\Theta(\mathsf{H}).$
  2. 2.
    $\Theta(\mathsf{G}^{i}\boxtimes\mathsf{H}^{j})\,\Theta(\mathsf{G})^{j}\,\Theta(\mathsf{H})^{i}$ $=\Theta(\mathsf{G}^{i}\boxtimes\mathsf{H}^{j})\,\Theta(\mathsf{G}^{j})\,\Theta(\mathsf{H}^{i})$
    $\leq\Theta(\mathsf{G}^{i+j}\boxtimes\mathsf{H}^{i+j})$
    $=\Theta(\mathsf{G}\boxtimes\mathsf{H})^{i+j}$
    $=\Theta(\mathsf{G})^{i+j}\,\Theta(\mathsf{H})^{i+j},$
    thus
    $\Theta(\mathsf{G}^{i}\boxtimes\mathsf{H}^{j})\leq\Theta(\mathsf{G})^{i}\,\Theta(\mathsf{H})^{j}.$ (35)
    Using property (35), for all $k\in\mathbb{N}$, we get
    $=\left(\Theta(\mathsf{G})+\Theta(\mathsf{H})\right)^{k}.$
    Finally, letting $k\to\infty$ gives
    $\Theta(\mathsf{G}+\mathsf{H})\leq\Theta(\mathsf{G})+\Theta(\mathsf{H}),$
    and by Theorem 2.11,
    $\Theta(\mathsf{G}+\mathsf{H})=\Theta(\mathsf{G})+\Theta(\mathsf{H}).$
@@
2.3 The Lovász function of graphs
Before we present the Lovász $\vartheta$-function of graphs, we define an orthogonal representation of a graph [11].

### v2

@@ Abstract
We derive exact values and new bounds for the Shannon capacity of two families of graphs: the $q$-Kneser graphs and the tadpole graphs. We also construct a countably infinite family of connected graphs whose Shannon capacity is not attained by the independence number of any finite strong power. Building on recent work of Schrijver, we establish sufficient conditions under which the Shannon capacity of a polynomial in graphs, formed via disjoint unions and strong products, equals the corresponding polynomial of the individual capacities, thereby reducing the evaluation of such capacities to that of their components. Finally, we prove an inequality relating the Shannon capacities of the strong product of graphs and their disjoint union, which yields streamlined proofs of several known bounds. In addition to contributing to the computation of the Shannon capacity of graphs, this paper is intended to serve as an accessible entry point to those wishing to work in this area.
Keywords. Shannon capacity of graphs, zero-error information theory, strong graph product; graph polynomials, independence number, Kneser graphs, tadpole graphs.
2020 Mathematics Subject Classification. 94A15 (primary), 05C35, 05C50, 05C69, 05C76 (secondary).
@@ 1 Introduction
The Shannon capacity of a graph, introduced by Claude Shannon in 1956 [1], is a central concept bridging zero-error information theory and graph theory. It characterizes the maximum rate of zero-error communication over a noisy channel when the channel is modeled by a graph whose vertices represent the input symbols, and two vertices are adjacent if and only if the corresponding symbols can be confused by the channel with positive probability. This notion establishes a fundamental link between zero-error problems in information theory and graph theory, as surveyed extensively in [2, 3, 4, 5, 6].
@@
While exploring properties of the Shannon capacity, it is worth keeping in mind the more general concept of the asymptotic spectrum of graphs, introduced by Zuiddam [22, 23, 21, 20], which delineates a space of graph parameters that remain invariant under graph isomorphisms. This space is characterized by the following unique properties: additivity under disjoint union of graphs, multiplicativity under strong product of graphs, normalization for a simple graph with a single vertex, and monotonicity under graph complement homomorphisms. Building upon Strassen’s theory of the asymptotic spectra [24], a novel dual characterization of the Shannon capacity of a graph is derived in [22], expressed as the minimum over the elements of its asymptotic spectrum. By confirming that various graph invariants, including the Lovász $\vartheta$-function [11] and the fractional Haemers’ bound [14], are elements of the asymptotic spectrum of a graph (spectral points), it can be deduced that these elements indeed serve as upper bounds on the Shannon capacity of a graph. For further exploration, the comprehensive paper by Wigderson and Zuiddam [23] provides a survey on Strassen’s theory of the asymptotic spectra and its application areas, including the Shannon capacity of graphs.
Several further properties of the Shannon capacity have been studied. For example, the recent work [25] analyzes the Shannon capacity of the distance-$k$ power of a graph using tools from spectral graph theory and linear optimization methods. Despite this progress, many fundamental questions remain open. Among the most basic is the determination of the Shannon capacity of odd cycle graphs of length greater than 5 [11, 27, 20, 26].
@@
  • Section 5 determines sufficient conditions for the unattainability of the Shannon capacity by the independence number of any finite strong power of a graph. It first presents in an alternative streamlined way an approach by Guo and Watanabe [10]. It then introduces two other original approaches. One of the novelties in this section is the construction of a countably infinite family of connected graphs whose capacity is unattainable by any finite strong power of these graphs.
  • Section 6 derives the Shannon capacity of the $q$-Kneser graphs, using a generalized result of Erdős-Ko-Rado [29], and a result on the spectrum of the $q$-Kneser graphs [30]. This broadens the class of graphs for which the Shannon capacity is explicitly known.
  • Section 7 introduces a new inequality relating the Shannon capacity of the strong power of graphs to that of their disjoint union, and identifies several conditions under which equality holds. As an illustration of the applicability of this inequality, this section also presents an alternative shorter proof of a lower bound for the disjoint union of a graph with its complement, originally obtained by N. Alon [7], along with analogous bounds on the Lovász $\vartheta$-function of graphs.
@@ 2.1.1 Terminology
  • A walk is a sequence of vertices in a graph $\mathsf{G}$, where every two consecutive vertices in the sequence are adjacent in $\mathsf{G}$.
  • A path is a walk with no repeated vertices.
  • A cycle $\operatorname{\mathsf{C}}$ is obtained from a path $\operatorname{\mathsf{P}}$ by adding an edge connecting the two endpoints of $\operatorname{\mathsf{P}}$ (i.e., an edge connecting the two degree-1 vertices of $\operatorname{\mathsf{P}}$).
@@
  • A path on $n$ vertices is denoted by $\operatorname{\mathsf{P}}_{n}$, and its size is equal to n1n-1.
  • A cycle on $n$ vertices is called an $n$-cycle, and it is denoted by $\operatorname{\mathsf{C}}_{n}$ with an integer $n\geq 3$. The order and size of $\operatorname{\mathsf{C}}_{n}$ are both equal to $n$.
@@
  • An $r$-partite graph is a graph whose vertex set is a disjoint union of $r$ subsets such that no two vertices in the same subset are adjacent. If $r=2$, then $\mathsf{G}$ is a bipartite graph (e.g., $\operatorname{\mathsf{C}}_{n}$ is a bipartite graph if and only if $n\geq 4$ is even).
@@
Throughout this paper, graphs are finite, simple, and undirected. We also use the standard notation $n]\triangleq\{1,\ldots,n\}$ for all $n\in\mathbb{N}$.
@@ 2.1.2 Graph operations
Define the $k$-fold strong power of $\mathsf{G}$ as the strong product of $k$ copies of $\mathsf{G}$:
Throughout this paper, we also use disjoint union of graphs as the addition operation on graphs.
@@
Let $\mathsf{G}$ and $\mathsf{H}$ be two simple graphs. The disjoint union $\mathsf{G}+\mathsf{H}$ is a graph whose vertices set is $\operatorname{\mathsf{V}}(\mathsf{G})\cup\operatorname{\mathsf{V}}(\mathsf{H})$, and the edges set is $\operatorname{\mathsf{E}}(\mathsf{G})\cup\operatorname{\mathsf{E}}(\mathsf{H})$. Also, for $m\in\mathbb{N}$, let
$m\mathsf{G}\triangleq\mathsf{G}+\ldots+\mathsf{G},$ (2)
which is defined as the disjoint union of $m$ copies of $\mathsf{G}$.
@@ 2.1.4 Fractional invariants of graphs
To properly define four fractional invariants of a graph, the following sets of functions are first introduced.
@@
Let $\mathsf{G}$ be a simple graph. Define four sets of functions as follows:
  • $\mathcal{F}_{I}(\mathsf{G})$ is the set of non-negative functions, $f\colon\operatorname{\mathsf{V}}(\mathsf{G})\to\mathbb{R}$, such that for every $\mathcal{I}\in\operatorname{\mathcal{I}}(\mathsf{G})$,
@@
  • $\mathcal{F}_{C}(\mathsf{G})$ is the set of non-negative functions, $f\colon\operatorname{\mathsf{V}}(\mathsf{G})\to\mathbb{R}$, such that for every $\mathcal{C}\in\operatorname{\mathcal{C}}(\mathsf{G})$,
@@
  • $\mathcal{G}_{I}(\mathsf{G})$ is the set of non-negative functions, $g\colon\operatorname{\mathcal{I}}(\mathsf{G})\to\mathbb{R}$, such that for every $v\in\operatorname{\mathsf{V}}(\mathsf{G})$,
    $\sum_{\mathcal{I}\in\operatorname{\mathcal{I}}(\mathsf{G}):\,v\in\mathcal{I}}g(\mathcal{I})\geq 1,$ (11)
    where the summation on the left-hand side of (11) is over all independent sets in $\mathsf{G}$ that include the vertex $v$.
  • $\mathcal{G}_{C}(\mathsf{G})$ is the set of non-negative functions, $g\colon\operatorname{\mathcal{C}}(\mathsf{G})\to\mathbb{R}$, such that for every $v\in\operatorname{\mathsf{V}}(\mathsf{G})$,
    $\sum_{\mathcal{C}\in\operatorname{\mathcal{C}}(\mathsf{G}):\,v\in\mathcal{C}}g(\mathcal{C})\geq 1,$ (12)
    where the summation on the left-hand side of (12) is over all cliques in $\mathsf{G}$ that include the vertex $v$.
Four basic fractional invariants of graphs are next defined by using linear programming.
@@
The next theorem follows from strong duality in linear programming.
@@
[37] Let $\mathsf{G}$ be a simple graph. Then
@@
From now on, we will primarily use the fractional independence number and the fractional chromatic number. When computing these quantities, we will use both of their equivalent linear programming formulations.
Another useful property, which is immediate from Definition 2.13, is given next.
@@
[37] For a simple graph $\mathsf{G}$, the following holds:
@@
Some properties of the fractional independence number are next presented.
@@ 2.1.6 Some structured families of graphs
A corollary by Hales [32] regarding the connection between the graph universality and fractional independence number is presented next.
@@
A graph $\mathsf{G}$ is strongly regular with parameters $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ if it satisfies
@@
  • Two distinct vertices $a,b\in\mathbb{F}_{q}$ are adjacent if aba-b is a square in $\mathbb{F}_{q}$.
@@
[33] Let $q\equiv 1\text{ mod 4}$ be a prime power, and let $\mathsf{G}=P(q)$ be a Paley graph. Then,
@@ 2.2 The Shannon capacity of graphs
The concept of the Shannon capacity of a graph $\mathsf{G}$ was introduced in [1] to consider the largest information rate that can be achieved with zero-error communication. A discrete memoryless channel (DMC) with finite input and output alphabets is specified by a finite input set $\mathcal{X}$, a finite output set $\mathcal{Y}$, and a non-empty fan-out set $\mathcal{S}_{x}\triangleq\bigl\{y\in\mathcal{Y}:\,p_{Y|X}(y|x)>0\bigr\}$ for each $x\in\mathcal{X}$. In each channel use, the sender transmits an input $x\in\mathcal{X}$, and the receiver observes an arbitrary output in the corresponding fan-out set $\mathcal{S}_{x}$. It is possible to represent the channel by a confusion graph $\mathsf{G}$, which is defined as follows:
  • $\operatorname{\mathsf{V}}(\mathsf{G})=\mathcal{X}$ represents the set of symbols of the channel’s input alphabet.
  • $\operatorname{\mathsf{E}}(\mathsf{G})$ is the edge set of $\mathsf{G}$, where two distinct vertices are adjacent if the corresponding two input symbols from $\mathcal{X}$ are not distinguishable by the channel, i.e., they can produce an identical output symbol with positive probability. In other words, two vertices in $\mathsf{G}$ are adjacent if the fan-out sets of the corresponding input symbols intersect, so
@@
The largest cardinality of a set of input symbols that can be communicated without error in a single use of the channel equals the independence number $\alpha(\mathsf{G})$. Indeed, in this single-use setting, the sender and the receiver agree in advance on an independent set $\mathcal{I}$ of a maximum size $\alpha(\mathsf{G})$, the sender transmits only inputs in $\mathcal{I}$, every received output is in the fan-out set of exactly one input in $\mathcal{I}$, and the receiver can correctly determine the transmitted input.
Consider the transmission of $k$-length strings over a channel, where the channel is used $k\geq 1$ times. The sender transmits a sequence $x_{1},\ldots,x_{k}$, and the receiver observes an output sequence $y_{1},\ldots,y_{k}$, where $y_{i}\in\mathcal{S}_{x_{i}}$ for all $i\in[k]$. In this setup, $k$ uses of the channel are viewed as a single use of a super-channel whose input alphabet is $\mathcal{X}^{k}$, its output alphabet is $\mathcal{Y}^{k}$, and the fan-out set of $(x_{1},\ldots,x_{k})\in\mathcal{X}^{k}$ is given by the cartesian product $\mathcal{S}_{x_{1}}\times\ldots\times\mathcal{S}_{x_{k}}$. Note that two distinct sequences $(x_{1},\ldots,x_{k}),(x_{1}^{\prime},\ldots,x_{k}^{\prime})\in\mathcal{X}^{k}$ are distinguishable by the channel if and only if
$S_{x_{i}}\cap S_{x_{i}^{\prime}}=\varnothing$ (25)
for some index $i\in[k]$. Thus, it is possible to represent the super-channel by the $k$-th confusion graph that is defined as the $k$-fold strong power of $\mathsf{G}$. Using the $k$-th confusion graph, the largest amount of information that can be sent by $k$ uses of the channel with error-free communication is given by the independence number of the $k$-th confusion graph, i.e., it is equal to $\alpha(\mathsf{G}^{k})$. The maximum information rate per symbol that is achieved by $k$ uses of the channel, and with zero-error communication, is equal to
$\frac{1}{k}\log\alpha(\mathsf{G}^{k})=\log\sqrt[k]{\alpha(\mathsf{G}^{k})},\quad k\in\mathbb{N}.$ (26)
Finally, by omitting the logarithm (as a monotonically increasing function) and supremizing over $k\in\mathbb{N}$, the Shannon capacity of the graph $\mathsf{G}$ is defined as follows.
@@
$\Theta(\mathsf{G})$ $\triangleq\sup_{k\in\mathbb{N}}\sqrt[k]{\alpha(\mathsf{G}^{k})}$ (27)
$=\lim_{k\to\infty}\sqrt[k]{\alpha(\mathsf{G}^{k})},$ (28)
where equality (28) holds by inequality (5) and Fekete’s Lemma.
We next provide several properties of the Shannon capacity, which are used throughout this paper.
@@
Let $\mathsf{G}$ be a simple graph and let $m\in\mathbb{N}$. Then,
$\Theta(\mathsf{G}^{m})=\Theta(\mathsf{G})^{m}.$ (29)
Proof.
Equality (29) holds by (28). ∎
@@
Let $\mathsf{G}_{1}$ and $\mathsf{G}_{2}$ be simple graphs. Then,
@@
[1] Let $\mathsf{G}_{1}$ and $\mathsf{G}_{2}$ be simple graphs. Then,
@@
Proof.
An elegant proof of Shannon’s inequality (31) is presented in [28]. ∎
In light of Definition 2.5, if $\mathsf{G}$ and $\mathsf{H}$ are the confusion graphs of two channels, then their disjoint union represents the sum of the channels corresponding to the situation where either one of the two channels may be used, a new choice being made for each transmitted symbol. The Shannon capacity of the disjoint union of graphs may be strictly larger than the sum of their capacities [7], disproving Shannon’s conjecture in the following strong sense.
Theorem 2.11 (The Shannon capacity of disjoint union of graphs).
[7, Theorem 1.1] For every $k\in\mathbb{N}$, there is a graph $\mathsf{G}$ so that the Shannon capacity of the graph and that of its complement $\overline{\mathsf{G}}$ satisfy $\Theta(\mathsf{G})\leq k$ and $\Theta(\overline{\mathsf{G}})\leq k$, whereas
$\Theta(\mathsf{G}+\overline{\mathsf{G}})\geq k^{(1+o(1))\,\frac{\ln k}{8\ln\ln k}},$ (32)
and the $o(1)$-term tends to zero as we let $k$ tend to infinity.
The following known result is central in this paper, so its short and nice proof from [28] is included here.
@@
[23, 28] Let $\mathsf{G}$ and $\mathsf{H}$ be simple graphs. Then
@@
See Appendix A. ∎
Theorem 2.13.
Let $\mathsf{G}$ be a simple graph and let $m\in\mathbb{N}$. Then,
$\Theta(m\mathsf{G})=m\,\Theta(\mathsf{G}).$ (34)
Proof.
Equality (34) follows from Theorem 2.12 by relying on equality (29). ∎
@@
Proof.
This holds by the definitions of induced and spanning subgraphs, and the definition in (27). ∎
2.3 The Lovász $\vartheta$-function of graphs
Before we present the Lovász $\vartheta$-function of graphs, we define an orthogonal representation of a graph [11] (see also [12, Chapter 11]).